# LeafNet: A Large-Scale Dataset and Comprehensive Benchmark for Foundational Vision-Language Understanding of Plant Diseases

### v1

@@ Abstract
Foundation models and vision-language pre-training have significantly advanced Vision–Language Models (VLMs), enabling multimodal processing of visual and linguistic data. However, their application in domain-specific agricultural tasks, such as plant pathology, remains limited due to the lack of large-scale, comprehensive multimodal image–text datasets and benchmarks. To address this gap, we introduce LeafNet, a comprehensive multimodal dataset, and LeafBench, a visual question-answering benchmark developed to systematically evaluate the capabilities of VLMs in understanding plant diseases. The dataset comprises 186,000 leaf digital images spanning 97 disease classes, paired with comprehensive expert-annotated metadata, generating 13,950 question-answer pairs spanning six critical agricultural tasks: crop species identification, healthy-diseased classification, disease identification, symptom identification, pathogen classification, and taxonomic name classification. The questions assess various aspects of plant pathology understanding, including visual symptom recognition, taxonomic relationships, and diagnostic reasoning. Benchmarking 12 state-of-the-art VLMs on our LeafBench dataset, including GPT-4o and Gemini 2.5 Pro, we reveal substantial disparity in their disease understanding capabilities. Closed-source models achieved up to 72% accuracy, whereas generic open-source VLMs struggle with near-chance performance. Our study shows performance varies markedly across tasks: binary healthy–diseased classification exceeds 90% accuracy, while fine-grained pathogen and species identification remains below 65%. Direct comparison between vision-only models and VLMs demonstrates the critical advantage of multimodal architectures: fine-tuned VLMs outperform traditional vision models by up to 27.76% on semantically demanding tasks, with domain-specific models like SCOLD achieving 99.15% disease identification accuracy and 94.92% symptom identification accuracy, confirming that integrating linguistic representations significantly enhances diagnostic precision. These findings highlight critical gaps in current VLMs for plant pathology applications and underscore the need for LeafBench as a rigorous framework for methodological advancement and progress evaluation toward reliable Artificial Intelligence (AI)-assisted plant disease diagnosis. Codes will be available at https://github.com/EnalisUs/LeafBench
@@
Keywords Vision-Language Models $\cdot$ Foundation Models $\cdot$ Plant Disease Diagnosis $\cdot$ Multimodal Dataset $\cdot$ Visual Question Answering $\cdot$ Benchmarking
@@ 1 Introduction
According to the 2025 report of the United Nations Food and Agriculture Organization (FAO), food insecurity and poverty remain critical, affecting more than 638 million people in 2024 due to growing demand and insufficient supply [1]. According to FAO, up to 40% of global crop yields are lost annually to pests and pathogens, resulting in economic losses exceeding USD 200 billion. This underscores the urgent need for more advanced data-driven methods to support informed decision-making in disease management and crop protection [2]. Digital agriculture plays a pivotal role in monitoring and managing agricultural pests and diseases through advanced data collection and analysis, thereby mitigating yield losses and enhancing global food security [3]. Recent studies have demonstrated the effectiveness of integrating digital technologies, including Internet of Things with Deep Learning (DL) [4], autonomous vehicle deployment driven by machine learning and computer vision [5], and the application of computer vision–based DL approaches [6] for precision disease management. Despite these advances, significant challenges remain due to environmental variability and the overlapping visual symptoms of different diseases, which hinder accurate identification. To overcome these limitations, our study introduces a large-scale digital image dataset and a language benchmark designed to support the development of robust VLMs, enabling not only image-based disease classification but also multimodal linguistic integration for zero-shot learning and Visual Question Answering (VQA).
@@ 3.2 LeafBench Detail
By associating each task with a predefined set of candidate labels, the question prompts explicitly guide the model toward discriminative reasoning rather than unconstrained text generation. This design enables reliable output parsing, facilitates automated accuracy computation, and ensures consistent, context-bounded inference across LeafNet’s diverse biological hierarchy. Importantly, all questions are designed to be "visual-dependent", such that correct answers cannot be derived from the text prompt alone and instead require analysis of the accompanying image.
@@ 3.3.2 Evaluation Tasks
Few-shot classification and transfer learning: Agricultural environments frequently encounter "long-tail" distributions where data for emerging pathogens or rare physiological disorders are scarce [50]. To simulate these real-world constraints, we employ a few-shot classification protocol. This task assesses the model’s ability to generalize to disease ontologies using limited supervision. We adopt an $N$-shot evaluation setting, where the model is trained on a randomly sampled support set containing $N$ examples per class, with $N\in\{16,32,64\}$. This logarithmic progression allows us to analyze the model’s learning curve and determine the minimal data requirement for effective disease detection.
@@ 4.4 Zero-Shot Visual Question Answering
Dataset difficulty and discriminative power: The LeafBench instruction set provides a significant challenge to off-the-shelf foundation models. As shown in Table 7, generic zero-shot baselines (e.g., CLIP, SigLIP2) frequently perform near the random chance lower bound ($\sim$24-30%) on fine-grained tasks such as SI and PC. This “generalization gap" confirms that LeafBench captures specialized pathological features that are underrepresented in general web-scale pre-training data. Conversely, the high performance of frontier proprietary models (GPT-4o, Gemini 2.5 Pro) and the domain-specialized SCOLD architecture (reaching 96.28% on Disease Identification) demonstrates that the dataset is solvable and contains distinct, learnable signal—establishing a meaningful expert upper bound.

### v2

@@ Abstract
Foundation models and vision-language pre-training have significantly advanced Vision–Language Models (VLMs), enabling multimodal processing of visual and linguistic data. However, their application in domain-specific agricultural tasks, such as plant pathology, remains limited due to the lack of large-scale, comprehensive multimodal image–text datasets and benchmarks. To address this gap, we introduce LeafNet, a comprehensive multimodal dataset, and LeafBench, a visual question-answering benchmark developed to systematically evaluate the capabilities of VLMs in understanding plant diseases. The dataset comprises 186,000 leaf digital images spanning 97 disease classes, paired with comprehensive expert-annotated metadata, generating 13,950 question-answer pairs spanning six critical agricultural tasks: crop species identification, healthy-diseased classification, disease identification, symptom identification, pathogen classification, and taxonomic name classification. The questions assess various aspects of plant pathology understanding, including visual symptom recognition, taxonomic relationships, and diagnostic reasoning. Benchmarking 12 state-of-the-art VLMs on our LeafBench dataset, including GPT-4o and Gemini 2.5 Pro, we reveal substantial disparity in their disease understanding capabilities. Closed-source models achieved up to 72% accuracy, whereas generic open-source VLMs struggle with near-chance performance. Our study shows performance varies markedly across tasks: binary healthy–diseased classification exceeds 90% accuracy, while fine-grained pathogen and species identification remains below 65%. Direct comparison between vision-only models and VLMs demonstrates the critical advantage of multimodal architectures: fine-tuned VLMs outperform traditional vision models by up to 27.76% on semantically demanding tasks, with domain-specific models like SCOLD achieving 99.15% disease identification accuracy and 94.92% symptom identification accuracy, confirming that integrating linguistic representations significantly enhances diagnostic precision. These findings highlight critical gaps in current VLMs for plant pathology applications and underscore the need for LeafBench as a rigorous framework for methodological advancement and progress evaluation toward reliable Artificial Intelligence (AI)-assisted plant disease diagnosis. Codes will be available at https://github.com/EnalisUs/LeafBench.
keywords:
@@
Vision-Language Models , Foundation Models , Plant Disease Diagnosis , Multimodal Dataset , Visual Question Answering , Benchmarking
[…]
[1]organization=School of Electrical Engineering, Korea University, city=Seoul, postcode=02841, country=South Korea
[…]
[2]organization=Department of Integrative Biology, The University of Texas at Austin, city=Austin, postcode=TX 78712, country=USA
[…]
[3]organization=Department of Software Engineering, FPT University, city=Cantho city, postcode=CT 90000, country=Vietnam
{graphicalabstract}
Figure 1: The LeafNet Curation and Benchmarking Pipeline. (A), Curation process overview. Metadata is synthesized from authoritative sources (NIH, NIFA) to map raw images to biological taxonomies, including species, disease, pathogenic agent, and symptom descriptions. (B), Expert verification and benchmarking. All metadata-image pairs undergo agricultural expert review to filter noisy samples. From this verified data, we construct LeafBench, a curated subset designed to benchmark LVLMs on image classification, few-shot learning, and VQA.
{highlights}
Introduce LeafBench, a comprehensive benchmarking framework for multiple Visual Question Answering (VQA) tasks in plant disease detection and characterization.
@@ 1 Introduction
According to the 2025 report of the United Nations Food and Agriculture Organization (FAO), food insecurity and poverty remain critical, affecting more than 638 million people in 2024 due to growing demand and insufficient supply [1]. According to FAO, up to 40% of global crop yields are lost annually to pests and pathogens, resulting in economic losses exceeding USD 200 billion. This underscores the urgent need for more advanced data-driven methods to support informed decision-making in disease management and crop protection [2]. Digital agriculture plays a pivotal role in monitoring and managing agricultural pests and diseases through advanced data collection and analysis, thereby mitigating yield losses and enhancing global food security [3]. Recent studies have demonstrated the effectiveness of integrating digital technologies, including Internet of Things with Deep Learning (DL) [4], autonomous vehicle deployment driven by machine learning and computer vision [5], and the application of computer vision–based DL approaches [6] for precision disease management. Despite these advances, significant challenges remain due to environmental variability and the overlapping visual symptoms of different diseases, which hinder accurate identification. To overcome these limitations, our study introduces a large-scale digital image dataset and a language benchmark designed to support the development of robust VLMs, enabling not only image-based disease classification but also multimodal linguistic integration for zero-shot learning and VQA.
@@ 3.2 LeafBench Detail
By associating each task with a predefined set of candidate labels, the question prompts explicitly guide the model toward discriminative reasoning rather than unconstrained text generation. This design enables reliable output parsing, facilitates automated accuracy computation, and ensures consistent, context-bounded inference across LeafNet’s diverse biological hierarchy. Importantly, all questions are designed to be ”visual-dependent”, such that correct answers cannot be derived from the text prompt alone and instead require analysis of the accompanying image.
@@ 3.3.2 Evaluation Tasks
Few-shot classification and transfer learning: Agricultural environments frequently encounter ”long-tail” distributions where data for emerging pathogens or rare physiological disorders are scarce [50]. To simulate these real-world constraints, we employ a few-shot classification protocol. This task assesses the model’s ability to generalize to disease ontologies using limited supervision. We adopt an $N$-shot evaluation setting, where the model is trained on a randomly sampled support set containing $N$ examples per class, with $N\in\{16,32,64\}$. This logarithmic progression allows us to analyze the model’s learning curve and determine the minimal data requirement for effective disease detection.
@@ 4.4 Zero-Shot Visual Question Answering
Dataset difficulty and discriminative power: The LeafBench instruction set provides a significant challenge to off-the-shelf foundation models. As shown in Table 7, generic zero-shot baselines (e.g., CLIP, SigLIP2) frequently perform near the random chance lower bound ($\sim$24-30%) on fine-grained tasks such as SI and PC. This “generalization gap” confirms that LeafBench captures specialized pathological features that are underrepresented in general web-scale pre-training data. Conversely, the high performance of frontier proprietary models (GPT-4o, Gemini 2.5 Pro) and the domain-specialized SCOLD architecture (reaching 96.28% on Disease Identification) demonstrates that the dataset is solvable and contains distinct, learnable signal—establishing a meaningful expert upper bound.